OpenStudy (anonymous):

log3 (x+5) = log3 (2x+13)

OpenStudy (oaktree):

Subtract the right hand side from the left and you have\[\log_{3}(x+5) - \log_{3}(2x+13) = 0 \]

OpenStudy (oaktree):

This simplifies to \[\log_{3}\frac{ x+5 }{ 2x+13 } = 0\]

OpenStudy (anonymous):

i got 5

OpenStudy (oaktree):

This means that \[\frac{ x+5 }{ 2x+13 } = 3^0 = 1\]

OpenStudy (anonymous):

since the base is the same .. Then cancel the log and solve for x x+2=2x+13 Solve for x now

OpenStudy (oaktree):

Now you can solve. You should get x=-8.

OpenStudy (anonymous):

@OakTree No need for that ...

OpenStudy (oaktree):

@Eyad That's true, but it helps to demonstrate the reasoning instead of just stating that something is true.

OpenStudy (anonymous):

oh wait, -8 is right, sryy

OpenStudy (anonymous):

ok just set those equations equal to each other as x+5=2x+13 and then pull the terms with x to one side and the numbers to the other side so.. x= -8 right.. and thats it :)